15-th Conference of the International Linear Algebra Society

April 25, 2008

Professor Thomas Milligan

University of Central Oklahoma Edmond, OK USA

Dear Professor Milligan:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “On Euclidean Squared Distance Matrices (with Chi-Kwong Li and Michael Trossett)” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Eunkyung Ahn

Kyungpook National University Daegu Korea

Dear Professor Ahn:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “An extended Lie-Trotter formula and its applications (with Sejong Kim and Yongdo Lim)” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Milica Andjelic

Center for Research on Optimization and Control Aveiro Portugal

Dear Professor Andjelic:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “One upper bound for the largest eigenvalue of the signless Laplacian (with Slobodan Simic)” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Marina Arav

Georgia State University Atlanta, GA USA

Dear Professor Arav:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “Sign Patterns That Require Almost Unique Rank (with Frank Hall, Zhongshan Li, Assefa Merid, Yubin Gao)” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Ali Armandnejad

Vali-asr University of Rafsanjan Rafsanjan Iran

Dear Professor Armandnejad:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “Right gw-majorization on M_{n,m}” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor José Barría

Santa Clara University Santa Clara U.S.A.

Dear Professor Barría:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “The strong closure of the similarity orbit for a class of pairs of finite rank operators” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Gerald Bourgeois

Faculté de Luminy Marseille France

Dear Professor Bourgeois:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “About the logarithm function over the matrices” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Rafael Bru

Univ. Politécnica Valencia Spain

Dear Professor Bru:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “On some classes of Hmatrices (with Ljiljana Cvetković, Vladimir Kostić and Francisco Pedroche)” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Miguel Carriegos

Universidad de León León Spain

Dear Professor Carriegos:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “Reachability of regular switched linear systems” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Wei Cheng

National University of Defense Technology Changsha P.R.China

Dear Professor Cheng:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “One Type of Inverse Eigenvalue Problems in Quaternionic Quantum Mechanics (with Liang-gui Feng)” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Wei Cheng

National University of Defense Technology Changsha P.R.China

Dear Professor Cheng:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “The Constrained Solutions of Quaternionic Matrix Equations (with Liang-gui Feng)” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Mikhail Cherepniov

Moscow State University Moscow Russia

Dear Professor Cherepniov:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “New block algorithm for solving large sparse linear systems over GF(2)” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Cristina Corral

Universidad Politécnica de Valencia Valencia Spain

Dear Professor Corral:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “On Schur complements of H-matrices (with R. Bru, I. Giménez, J. Mas)” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Liliana Costa

Centre for Research on Optimization and Control Aveiro Portugal

Dear Professor Costa:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “Acyclic Birkhoff Polytope (with C.M. da Fonseca and Enide Andrade Martins)” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Glória Cravo

University of Madeira and CELC Funchal Portugal

Dear Professor Cravo:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “Controllability of Matrices with Prescribed Blocks” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Slavisa Djordjevic

Benemérita Universidad Autónoma de Puebla Puebla México

Dear Professor Djordjevic:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “Manifold of proper elements (with S. Sánchez Perales)” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Marija Dodig

CELC, Universidade de Lisboa Lisbon Portugal

Dear Professor Dodig:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “Singular systems, state feedback problem” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Gregor Dolinar

Faculty of Electrical Engineering Ljubljana Slovenia

Dear Professor Dolinar:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “General preservers of quasicommutativity (with Bojan Kuzma)” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Esther Dopazo

Facultad de Informatica. Universidad Politécnica Madrid Boadilla del Monte, Madrid Spain

Dear Professor Dopazo:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “Further results on the representation of the Drazin inverse of a 2×2 block matrix (with M.F. Martínez-Serrano and N. Castro-González)” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Rosário Fernandes

Centro de Estruturas Lineares e Combinatórias (CELC) Lisboa Portugal

Dear Professor Fernandes:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “Rank partitions and covering numbers under small perturbations of an element” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Josep Ferrer

Universitat Politecnica de Catalunya Barcelona Spain

Dear Professor Ferrer:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “Geometric structure of the equivalence classes of a controllable pair (with A. Compta, and M. Peña)” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Carlos Fonseca

Department of Mathematics, University of Coimbra Coimbra Portugal

Dear Professor Fonseca:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “An inequality for the multiplicity of an eigenvalue” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Davoud Foroutannia

Vali-asr University of Rafsanjan Rafsanjan Iran

Dear Professor Foroutannia:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “Bounds for matrices on weighted sequence spaces” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Martin Frank

University of Kaiserslautern Kaiserslautern Germany

Dear Professor Frank:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “An iterative method for transport equations in radiotherapy (with Bruno Dubroca)” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Susana Furtado

Centro de Estruturas Lineares e Combinatórias da U. L Porto Portugal

Dear Professor Furtado:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “Order Invariant Spectral Properties for Several Matrices (with Charles Johnson)” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Shigeru Furuichi

Nihon University Tokyo Japan

Dear Professor Furuichi:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “On trace inequalities for products of matrices” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Lev Glebsky

Universidad Autónoma de San Luis Potosi San Luis Potosi México

Dear Professor Glebsky:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “On low rank perturbations of matrices (with Luis Manuel Rivera)” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Assaf Goldberger

Tel Aviv University Tel Aviv Israel

Dear Professor Goldberger:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “An upper bound on the characteristic polynomial of a nonnegative matrix leading to a proof of the Boyle–Handleman conjecture (with Michael Neumann)” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Maria Gouveia

Department of Mathematics, FCTUC Coimbra Portugal

Dear Professor Gouveia:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “On a singular Toeplitz pencil” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Sergey Grudsky

CINVESTAV del I.P.N. México México

Dear Professor Grudsky:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “Uniform boundedness of Toeplitz Matrices with variable coefficients” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Ibrahim Hanaish

7th October University Misurata Libya

Dear Professor Hanaish:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “Shrinkage Estimators for Estimation the Multivariate Normal Mean Vector under Degrees of Distrust (with Abdunnabi M. Ali Elbouzedi)” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Roger Horn

University of Utah Salt Lake City USA

Dear Professor Horn:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “A Canonical Form for Quasi-Real Normal Matrices Under Real Orthogonal Similarity (with Geoffrey R. Goodson)” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Dirk Huylebrouck

Dep Architecture Sint-Lucas Brussels Belgium

Dear Professor Huylebrouck:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “Applications of generalized inverses in art” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Bokhee Im

Chonnam National University Gwangju Korea(Rep. of)

Dear Professor Im:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “Representations of trilinear products in Comtrans algebras (with Jonathan D. H. Smith)” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Hussein Kaibah

7th October University Misurata Libya

Dear Professor Kaibah:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “Asymptotic Behavior of Solutions of Stochastic Equations and Applications in Statistical Parameter Estimation” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Peter Kortesi

University of Miskolc Miskolc Hungary

Dear Professor Kortesi:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “Using Linear Algebra in Teaching Hamilton Quaternions and Graphs” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Gue Myung Lee

Pukyong National University Busan Korea

Dear Professor Lee:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “Complexity Analysis of the Primal-Dual Interior Point Method for Second-order Cone Optimization Problem (with Bo Kyung Choi)” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Silvia Machado

Pontificia Universidade Catolica de São Paulo São Paulo Brasil

Dear Professor Machado:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “GPEA’s researches about the meta resources in teaching and learning the notion of basis of a vector space (with Bianchini, B.L. and Maranhão, M.C.S.A.)” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor John Maroulas

Dept. of Mathematics, Nat. Technical University Athens Greece

Dear Professor Maroulas:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “Dilation of numerical ranges of normal matrices (with Maria Adam)” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Janko Marovt

Institute of Mathematics, Physics and Mechanics Ljubljana Slovenia

Dear Professor Marovt:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “Homomorphisms of matrix semigroups over division rings from dimension two to three (with Gregor Dolinar)” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Maria-da-Graça Marques

University of Algarve and CELC Faro Portugal

Dear Professor Marques:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “Some conditions for the commutativity of matrix patterns (with C.R. Johnson)” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Enide Martins

Centre for Research on Optimization and Control (CEOC) Aveiro Portugal

Dear Professor Martins:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “On the spectra of some graphs like weighted rooted trees (with Rosário Fernandes and Helena Gomes)” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Rana Mikkelson

Iowa State University Ames, IA United States of America

Dear Professor Mikkelson:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “Minimum Rank of Graphs with Loops” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Lon Mitchell

Virginia Commonwealth University Richmond United States

Dear Professor Mitchell:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “Orthogonal Removal of Vertices and Minimum Semidefinite Rank (with Sivaram Narayan)” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Sivaram Narayan

Central Michigan University Mount Pleasant, Michigan 48859 USA

Dear Professor Narayan:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “Linearly Independent Vertices and Minimum Semidefinite Rank” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Vadim Olshevsky

University of Connecticut Storrs USA

Dear Professor Olshevsky:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “Can One Invert a Matrix via Graph Manipulations? (with Tom Bella and Pavel Zhlobich)” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Alejandro Palma

Instituto de Física (BUAP) Puebla México

Dear Professor Palma:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “Solution of the linear timedependent potential by using a solvable Lie algebra^{∗} (with M. Villa, and L. Sandoval)” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Pedro Patricio

Departamento de Matemática, Universidade do Minho Braga Portugal

Dear Professor Patricio:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “Some additive results on Drazin Inverses (with R.E. Hartwig)” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor George Poole

East Tennessee State University Johnson City USA

Dear Professor Poole:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “Whatever Happened to Rook’s Pivoting?” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Volodymyr Prokip

Institute of Appl.Problem for Mech.and Math. NAS of Ukr L’viv Ukraine

Dear Professor Prokip:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “On the problem of diagonalizability of matrices over a principal ideal domain” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Vladimir Protasov

Moscow State University Moscow Russia

Dear Professor Protasov:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “p-radii of linear operators and equations of self-similarity” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Rosa E. Pruneda

University of Castilla-La Mancha Ciudad Real Spain

Dear Professor Pruneda:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “Complete Orthogonal Decomposition Compared with Direct Projection Methods (with Beatriz Lacruz)” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Fanja Rakotondrajao

University of Antananarivo Antananarivo Madagascar

Dear Professor Rakotondrajao:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “Euler’s difference table and maximum permanents of (0, 1)-matrices” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Alicia Roca

Dpto. de Matemática Aplicada, Universidad Politécnica Valencia Spain

Dear Professor Roca:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “Pencils with Prescribed Constant Subpencils (with F.C. Silva)” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Siegfried M. Rump

Hamburg University of Technology Hamburg Germany

Dear Professor Rump:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “The ratio between the Toeplitz and the unstructured condition number (with H. Sekigawa)” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Maria Rosaria Russo

Department of Mathematics - University of Padua Padova Italy

Dear Professor Russo:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “On some general determinantal identities of Sylvester type (with Michela Redivo-Zaglia)” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Ahmed Salam

Université du Littoral-Côte d’Opale Calais France

Dear Professor Salam:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “A structure-preserving Arnoldi-like method for a class of structured matrices” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Elisa Schaeffer

Universidad Autónoma de Nuevo León San Nicolás de los Garza México

Dear Professor Schaeffer:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “Locally computable approximations of absorption times for graph clustering (with Pekka Orponen and Vanesa Avalos)” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Burkhard Schaffrin

Ohio State University Columbus, OH USA

Dear Professor Schaffrin:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “Total least-squares regularization of Tykhonov type and an ancient racetrack in Corinth (with Kyle Snow)” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Anatoly Sebeldin

University UGANC Guinea, Conakry Guinea

Dear Professor Sebeldin:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “Algorithm resolving problem of determination of finite cyclicgroup by its automorphism group (with V.K. Vildanov and A.L. Sylla)” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Morteza Seddighin

Indiana University East Richmod Indiana

Dear Professor Seddighin:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “Matrix Optimization in Statistics” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Peter Semrl

University of Ljubljana Ljubljana Slovenia

Dear Professor Semrl:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “Locally linearly dependent operators” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Hristo Sendov

The University of Western Ontario London Canada

Dear Professor Sendov:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “Spectral Manifolds (with A. Daniilidis, J. Malick, and A. Lewis)” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Mohammad Shahryari

Tabriz University, Tabriz, Iran Tabriz Iran

Dear Professor Shahryari:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “ℤ_{2}-graded symmetry classes of tensors” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor John Sinkovic

Technische Universiteit Eindhoven Eindhoven Netherlands

Dear Professor Sinkovic:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “An upper bound for the maximum nullity of a symmetric matrix whose graph is outerplanar” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Koratti Chengalrayan Sivakumar

Indian Institute of Technology Madras Chennai India

Dear Professor Sivakumar:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “Least Elements of Polyhedral Sets and Nonnegative Generalized Inverses (with Debashisha Mishra)” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Klemen Sivic

Institute of Mathematics, Physics and Mechanics Ljubljana Slovenia

Dear Professor Sivic:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “On varieties of commuting triples” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Graça Soares

University of Tras-os-Montes and Alto Douro Vila Real Portugal

Dear Professor Soares:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “Inequalities on an indefinite inner product space (with N. Bebiano et al.)” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Marko Stosic

Instituto de Sistemas e Robotica, IST Lisbon Portugal

Dear Professor Stosic:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “On Generalized Procrustes Problem (with João Xavier)” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Mitsuru Uchiyama

Shimane University, Professor Matsue, Shimane Japan

Dear Professor Uchiyama:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “A New Majorization between functions” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Frank Uhlig

Mathematics, Auburn University Auburn, AL 36849 USA

Dear Professor Uhlig:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “Convex and Non-convex Optimization Problems for the Field of Values of a Matrix” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Ernesto Vallejo

Instituto de Matemáticas Morelia México

Dear Professor Vallejo:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “Additivity obstructions for integral matrices and pyramids (with Miguel Santoyo)” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Luis Vieira

Feup Porto Portugal

Dear Professor Vieira:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “Euclidean Jordan algebras and inequalities on the parameters and on the spectra of a strongly regular graph” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor James Weaver

University of West Florida Pensacola, FL 32514 USA

Dear Professor Weaver:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “Nonsingularity of Divisor Tournaments (with Rohan Hemasinha and Jeffrey L. Stuart)” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Piotr Wojciechowski

University of Texas at El Paso El Paso USA

Dear Professor Wojciechowski:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “Orderings of matrix algebras and their applications” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Iwona Wrobel

Warsaw University of Technology and Polish Academy of Sciences Warsaw Poland

Dear Professor Wrobel:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “The Gauss-Lucas theorem and the numerical range” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Pei Yuan Wu

National Chiao Tung University Hsinchu Taiwan

Dear Professor Wu:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “Numerical ranges of nilpotent operators (with Hwa-Long Gau)” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx